# Substitution for minimizing/maximizing a tropical linear (fractional) programming

### 27 Mar 2026 v1

@@ 2.3 Organization of the paper
In the Section 3 the necessary materials to solve the minimization problem mpr. The substitution method is described in Section 4.
@@ 3.1 Elementary results from interval arithmetic in dioids
And clearly by Result 3.1 the set $\{y:y\lesseqgtr A\otimes y\oplus B\otimes u\}$ admits a smallest (or greatest) element iff condition (15) is verified. Here, the condition (15) is equivalent to $\lim_{k\rightarrow+\infty}C^{\otimes(k)}=\boldsymbol{O}_{1,1}$ which is equivalent to $a^{\otimes(-1)}\otimes v_{n}<\mathbbm{1}$ because the matrix $C$ has only one elementary circuit $n\rightarrow n$ of weight $a^{\otimes(-1)}\otimes v_{n}$. To conclude we just have to note that by defining $\overline{v}^{\intercal}:=(u^{\intercal},\mbox{{\scriptsize$\mathbb{O}$}})$ the inequality (18a) is verified. $\Box$
@@ 4 The minimization problem mpr
  • Case $2$: $b^{[n]}\ngeq d^{[n]}$. In this case $h=+\infty$ is the only possible value for $h$ because $h>\mbox{{\scriptsize$\mathbb{O}$}}$ and $\min(z)=+\infty$.
@@
Assuming that $I_{\geq}(x_{j},i)$ is valid we define the following function:
@@ 4.2.2 Choosing a variable for a new substitution after $k$ substitutions in mpr
Let $\mathbb{T}$ and $\mathbb{T}^{\prime}$ denoting either $\mathbb{B}$ or $\mathbb{U}$, respectively. The $2$-tuple $(\mathbb{T},\mathbb{T}^{\prime})$ is defined as a function of the set $\mathbb{I}$ according to $\preceq_{fct}$ as follows (priority to the cost function):
@@
The equivalence (45) comes from the properties of the familly of intervals we are considering in this paper. We have a total order (see Proposition 3.1) and the Min of a set of intervals coincides with the intersection of the same set of intervals (see (13)). The proof is achieved by noticing that $\forall(i,j)\in\textsf{argMin}(\tau)\cap\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I})$ the variables $x_{j}$ can be exchanged. And thus one have to take the Min. $\Box$
@@
Or Case $0.2$: $\mathbbm{x}\neq\emptyset$ and $I_{\leq}(\mathbbm{x})=\emptyset$ and $I_{\geq}(\mathbbm{x})=\emptyset$ and the condition (37) of Theorem 4.2 does not holds. The substitution process stops because no valid inequalities can be generated. No minimum reached.
@@
The case is Case $2.2$: $\mathbbm{x}\neq\emptyset$ and $\exists x_{j}\in\mathbbm{x}$ such that: $I_{\geq}(x_{j})\neq\emptyset$. Let us distinguish the following subcases which are based on the form of the cost function:
@@ 4.2.3 The new characteristic elements of the mpr after substitution
$\forall j:t^{k\rightarrow k+1}_{j,.}:=\left\{\begin{array}[]{lc}e_{j}\mbox{ if $j\neq j^{*}$}\\ (\mbox{{\scriptsize$\mathbb{O}$}},v^{\intercal}_{i^{*}j^{*}},r_{i^{*}j^{*}})&\mbox{ if $j=j^{*}$.}\end{array}\right$. (52)
Where $e_{j}$ denotes the $j$th $n+2$-dimensional row vector of the $(n+2)\times(n+2)$-identity matrix $\boldsymbol{I}_{n+2}$.
@@ 6.1 A minimum which is equal to $\infty$
Constraint $l_{2}$ implies that $h=+\infty$ is the only possible value for $h$, where $h$ is assumed to be $\mbox{{\scriptsize$\mathbb{O}$}}$. Then, by back substitution we have: $z=2\otimes+\infty=+\infty$ and $x_{1}=2\otimes+\infty=+\infty$. And $\infty$ is element of the set $\mathcal{P}(A,b,C,d)\cup\{\boldsymbol{\pm\infty}\}$.
@@ 6.2 Minimization Example
Thus, $(i^{*},j^{*})=(4,1)$ and the $4\times 4$-transition matrix $T^{0\rightarrow 1}$ is defined by:
@@ 6.3 Fractional Example
By Theorem 4.3: $(\mathbb{T},\mathbb{T}^{\prime})=(\mathbb{B},\mathbb{B})$. So, we have: $(i^{*},j^{*})=(2,2)$ from the previous array. The new cost function is $\textsf{cost}^{[2]}(y,t,h)=1\otimes t\oplus 4^{\otimes(-1)}\otimes h$.
And we define the $6\times 6$-transition matrix $T^{1\rightarrow 2}$ which corresponds to $y_{2}=1\otimes t\oplus 4^{\otimes(-1)}\otimes h$ as follows:
@@
Applying Theorem 4.3: $(\mathbb{T},\mathbb{T}^{\prime})=(\mathbb{B},\mathbb{B})$. We compute the interval $\tau$ and we have: $\tau=[4^{\otimes(-1)}\otimes\mu,+\infty]$. And $\textsf{argMin}(\tau)=\{(1,t),(4,t)\}$ is not a singleton. So, we have to compute $\tau^{\prime}$, Theorem 4.3. And $\tau^{\prime}=[5^{\otimes(-1)}\otimes\mu,+\infty]$. Then, we take $(i^{*},j^{*})\in\textsf{argMin}(\tau^{\prime})=\{(4,t)\}$. Thus, $\textsf{cost}^{[3]}(y,t,h)=3^{\otimes(-1)}\otimes y_{3}\oplus 4^{\otimes(-1)}\otimes h$. And we apply the results of § 4.2.3. Then, the set of linear equalities is now:
@@
And we define the $6\times 6$-transition matrix $T^{1\rightarrow 2}$ as follows:
@@
And we have the solution of the mpr obtained by obvious back substitution in $\mathcal{L}^{[4]}$:
@@ 7 Conclusion
Now we have to add the time complexity of the resolution of the triangular system of equalities $\mathcal{L}^{[n]}$ which is $\mathcal{O}(n)$ because we do not need to complute a pseudo inverse of a matrix by residuation theory (see eg. [3]).
@@
Let us recall that to the best of our knowledge this substitution method is the first strongly polynomial method for such problem. All other proposed schemes of resolution are pseudo-polynomial (see [5], [10], [11]). Thus, because $(\max,+)$-linear programming is now proved to be strongly polynomial then the mean payoff games problem is also strongly polynomial (see eg. [10]). Let us remark that this result was already proved in [19] when solving the max-atom problem (MAP). The important consequence of this result is that other six PTIME equivalent problems to MAP which were known to be in NP $\cap$ co-NP (see eg [1], [4], [7], [8], [9], [12], [16], [17], [18], [21]) are also strongly polynonial. We list them hereafter. P1: Looking for non trivial solutions of a tropical cone, P2: Computation of a tropical rank of a matrix, P3: Computation of optimal strategies in parity games (typically: Mean Payoff Games), P4: Scheduling with and/or precedence constraints, P5: Shortest path problem in hypergraph, P6: Model checking and $\mu$-calculus.
@@ A.2 Choosing a variable for a new substitution after $k$ substitutions in MPR
$A_{[|1,m|]h}^{+[k]}\leq A_{[|1,m|]h}^{-[k]}.$ (105)
@@
Case 0: $\max(z)=\mbox{{\scriptsize$\mathbb{O}$}}$ or no maximum. Case $0.1$: $\mathbbm{x}=\emptyset$ and the condition (105) of Theorem A.2 does not holds. $\mathbbm{x}=\emptyset$ means that $k=n$ (no more variable to be substituted) and $\textsf{cost}^{[n]}(x,h)=c_{h}^{[n]}\otimes h$. The condition (105) of Theorem A.2 does not holds means $A_{[|1,m|]h}^{+[n]}\nleq A_{[|1,m|]h}^{-[n]}$ then $h=\mbox{{\scriptsize$\mathbb{O}$}}$ is the only possible value for $h$ and by backward substitution we have $z=\mbox{{\scriptsize$\mathbb{O}$}}$, $x=\boldsymbol{-\infty}$.
Or Case $0.2$: $\mathbbm{x}\neq\emptyset$ and $I_{\leq}(\mathbbm{x})=\emptyset$ and $I_{\geq}(\mathbbm{x})=\emptyset$ and the condition (105) of Theorem A.2 does not holds. The substitution process stops because no valid inequalities can be generated. No maximum.
Case $1$: switching case, cf. § 2.4. $\mathbbm{x}^{+}=\emptyset$ and $\textsf{cost}^{[k]}(x,h)=c_{h}^{[k]}\otimes h$ and the condition (105) of Theorem A.2 does not holds and $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}\neq\emptyset$ and $I_{\leq}(\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}})=\emptyset$ and $I_{\geq}(\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}})\neq\emptyset$. The substitution must switch to the following minimizing problem:
@@
The case is Case $2.2$: $\mathbbm{x}\neq\emptyset$ and $\exists x_{j}\in\mathbbm{x}$: $I_{\leq}(x_{j})\neq\emptyset$. Let us distinguish the following subcases which are based on the form of the cost function:
@@
Case $2.2.2$: otherwise, the condition (105) of Theorem A.2 does not hold. Based on the hierarchy $\preceq_{\overline{var}}$ (see (101)) between the remaining variables let us define the set of the dominating variables $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}:=\{x_{j}\in\mathbbm{x}:\;I_{\geq}(x_{j})\neq\emptyset\mbox{ and }I_{\leq}(x_{j})\neq\emptyset\}$. And the set of inequalities:
@@ A.3 The new characteristic elements of the MPR after substitution
$\forall j:t^{k\rightarrow k+1}_{j,.}:=\left\{\begin{array}[]{lc}e_{j}\mbox{ if $j\neq j^{*}$}\\ (\mbox{{\scriptsize$\mathbb{O}$}},v^{\intercal}_{i^{*}j^{*}},r_{i^{*}j^{*}})&\mbox{ if $j=j^{*}$.}\end{array}\right$. (120)
Where $e_{j}$ denotes the $j$th $n+2$-dimensional row vector of the $(n+2)\times(n+2)$-identity matrix $\boldsymbol{I}_{n+2}$.
@@ A.4 A numerical example
The new characteristic elements are listed below.
@@
$\mathbbm{x}=\{x_{1}\}\neq\emptyset$. It is easy to see that the MPR problem is bounded. We have $\mathbbm{x}^{+}=\{x_{1}\}$ and $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}=\emptyset$. The set of dominating variables is $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}:=\{x_{1}\}$. The condition (105) of Theorem A.2 is not verified. Thus, Case $2.2.2$ applies. We have (see (117)): $\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D)=I_{\leq}(\{x_{1}\})=\{(2,1),(3,1),(4,1)\}$.
@@
The new characteristic elments of the problem are listed below.
@@
Clearly, $A_{[|1,4|]h}^{+[2]}=\boldsymbol{-\infty}\leq A_{[|1,4|]h}^{-[2]}=\boldsymbol{-\infty}$. And the condition (105) of Theorem A.2 is verified and the cost function $h\mapsto 5\otimes h$ is reached.
The solution of the MPR is obtained by a trivial back substitution:

### 02 Apr 2026 v2

@@ 2.3 Organization of the paper
Section 3 deals with the necessary materials to solve the minimization problem mpr. The substitution method is described in Section 4.
@@ 3.1 Elementary results from interval arithmetic in dioids
And clearly by Result 3.1 the set $\{y:y\lesseqgtr A\otimes y\oplus B\otimes u\}$ admits a smallest (or greatest) element iff condition (15) is verified. Here, the condition (15) is equivalent to $\lim_{k\rightarrow+\infty}C^{\otimes(k)}=\boldsymbol{O}_{1,1}$ which is equivalent to $a^{\otimes(-1)}\otimes v_{n}<\mathbbm{1}$ because the matrix $C$ has only one elementary circuit $n\rightarrow n$ of weight $a^{\otimes(-1)}\otimes v_{n}$. To conclude we just have to note that by defining $\overline{v}^{\intercal}:=(v_{1},\ldots,v_{n-1},\mbox{{\scriptsize$\mathbb{O}$}})$ the inequality (18a) is verified. $\Box$
@@ 4 The minimization problem mpr
  • Case $2$: $b^{[n]}\ngeq d^{[n]}$. In this case $h=+\infty$ is the only possible value for $h$ because $h>\mbox{{\scriptsize$\mathbb{O}$}}$ (see condition (6d) and $\min(z)=+\infty$.
@@
Assume that $I_{\geq}(x_{j},i)$ is valid. Then, by replacing $x_{j}$ by $f^{\geq}_{ij}(x,h)$ in the cost function $c^{\intercal}\otimes x\oplus c_{h}\otimes h$ we define the following function:
@@ 4.2.2 Choosing a variable for a new substitution after $k$ substitutions in mpr
Let $\mathbb{T}$ and $\mathbb{T}^{\prime}$ denoting either $\mathbb{B}$ or $\mathbb{U}$, respectively. The $2$-tuple $(\mathbb{T},\mathbb{T}^{\prime})$ is defined as a function of the set $\mathbb{I}$ according to $\preceq_{fct}$ as follows (recalling the priority to the cost function):
@@
The equivalence (45) comes from the properties of the familly of intervals we are considering in this paper. We have a total order (see Proposition 3.1) and the Min of a set of intervals coincides with the intersection of the same set of intervals (see (13)). The proof is achieved by noticing that $\forall(i,j)\in\textsf{argMin}(\tau)\cap\mathcal{T}(\mathbb{T},\mathbb{T}^{\prime},\mathbb{I})$ the variables $x_{j}$ can be exchanged. And thus we have to take the Min. $\Box$
@@
Case $0.2$: $\mathbbm{x}\neq\emptyset$ and $I_{\leq}(\mathbbm{x})=\emptyset$ and $I_{\geq}(\mathbbm{x})=\emptyset$ and the condition (37) of Theorem 4.2 does not holds. The substitution process stops because no valid inequalities can be generated. No minimum reached.
@@
The other case is Case $2.2$: $\mathbbm{x}\neq\emptyset$ and $\exists x_{j}\in\mathbbm{x}$ such that: $I_{\geq}(x_{j})\neq\emptyset$. Let us distinguish the following subcases which are based on the form of the cost function:
@@
Where: $I_{\geq}(\mathbbm{x}\cap D^{+\mbox{{\scriptsize$\mathbb{O}$}}}):=\{(i,j):\mbox{ $x_{j}\in\mathbbm{x}\cap D^{+\mbox{{\scriptsize$\mathbb{O}$}}}$ and $I_{\geq}(x_{j},i)$ is valid}\}$.
@@ 4.2.3 The new characteristic elements of the mpr after substitution
$\forall j:t^{k\rightarrow k+1}_{j,.}:=\left\{\begin{array}[]{lc}e^{\intercal}_{j}\mbox{ if $j\neq j^{*}$}\\ (\mbox{{\scriptsize$\mathbb{O}$}},v^{\intercal}_{i^{*}j^{*}},r_{i^{*}j^{*}})&\mbox{ if $j=j^{*}$.}\end{array}\right$. (52)
Where $e^{\intercal}_{j}$ denotes the $j$th $n+2$-dimensional row vector of the $(n+2)\times(n+2)$-identity matrix $\boldsymbol{I}_{n+2}$.
@@ 5 Linear fractional programming
And
$\mbox{$z$ and $h$ are not substituable}.$ (58e)
Finally, among the variables of the problem $z,y_{1},\ldots,y_{n},t$ we have:
[…]
@@ 6.1 A minimum which is equal to $\infty$
Constraint $l_{2}$ implies that $h=+\infty$ is the only possible value for $h$, where $h$ is assumed to be $\mbox{{\scriptsize$\mathbb{O}$}}$. Then, by backward substitution we have: $z=2\otimes+\infty=+\infty$ and $x_{1}=2\otimes+\infty=+\infty$. And $\infty$ is element of the set $\mathcal{P}(A,b,C,d)\cup\{\boldsymbol{\pm\infty}\}$.
@@ 6.2 Minimization Example
Thus, obviously $(i^{*},j^{*})=(4,1)$ which means that $x_{1}=2^{\otimes(-1)}\otimes h$, in the row $4$ of the linear system of inequalities (66b). We apply the results of § 4.2.3. Then, the set of linear equalities is now:
$\mathcal{L}^{[2]}=\{x_{2}=2\otimes h,x_{1}=2^{\otimes(-1)}\otimes h\}.$ (67)
The new cost function is:
$\textsf{cost}^{[2]}(x,h):=f_{z,32}(x,h)=h.$ (68)
The $4\times 4$-transition matrix $T^{1\rightarrow 2}$ is defined by:
@@ 6.3 Fractional Example
By Theorem 4.3: $(\mathbb{T},\mathbb{T}^{\prime})=(\mathbb{B},\mathbb{B})$. So, we have: $(i^{*},j^{*})=(2,2)$ from the previous array which means that $y_{2}=1\otimes t\oplus 4^{\otimes(-1)}\otimes h$.
@@
The new cost function is:
$\textsf{cost}^{[2]}(y,t,h)=1\otimes t\oplus 4^{\otimes(-1)}\otimes h.$ (80)
We define the $6\times 6$-transition matrix $T^{1\rightarrow 2}$ which corresponds to $y_{2}=1\otimes t\oplus 4^{\otimes(-1)}\otimes h$ as follows:
@@
Applying Theorem 4.3: $(\mathbb{T},\mathbb{T}^{\prime})=(\mathbb{B},\mathbb{B})$. We compute the interval $\tau$ and we have: $\tau=[4^{\otimes(-1)}\otimes\mu,+\infty]$. And $\textsf{argMin}(\tau)=\{(1,t),(4,t)\}$ is not a singleton. So, we have to compute $\tau^{\prime}$, Theorem 4.3. And $\tau^{\prime}=[5^{\otimes(-1)}\otimes\mu,+\infty]$. Then, we take $(i^{*},j^{*})\in\textsf{argMin}(\tau^{\prime})=\{(4,t)\}$ which means that $t=3^{\otimes(-1)}\otimes y_{3}\oplus 5^{\otimes(-1)}\otimes h$. We apply the results of § 4.2.3. Then, the set of linear equalities is now:
@@
And the new cost function is:
$\textsf{cost}^{[3]}(y,t,h)=3^{\otimes(-1)}\otimes y_{3}\oplus 4^{\otimes(-1)}\otimes h.$ (83)
We define the $6\times 6$-transition matrix $T^{1\rightarrow 2}$ as follows:
@@
And we have the solution of the mpr obtained by obvious backward substitution in $\mathcal{L}^{[4]}$:
@@ 7 Conclusion
Now we have to add the time complexity of the resolution of the triangular system of equalities $\mathcal{L}^{[n]}$ which is $\mathcal{O}(n)$ because we do not need to complute a pseudo inverse of a matrix by residuation theory (see eg. [3]). This is only backward substitution.
@@
Let us recall that to the best of our knowledge this substitution method is the first strongly polynomial method for such problem. All other proposed schemes of resolution are pseudo-polynomial (see [5], [10], [11]). Thus, because $(\max,+)$-linear programming is now proved to be strongly polynomial then the mean payoff games problem is also strongly polynomial (see eg. [10]). Let us remark that this result was already proved in [19] (english paper) when solving the max-atom problem (MAP). The important consequence of this result is that other six PTIME equivalent problems to MAP which were known to be in NP $\cap$ co-NP (see eg [1], [4], [7], [8], [9], [12], [16], [17], [18], [21]) are also strongly polynonial. We list them hereafter. P1: Looking for non trivial solutions of a tropical cone, P2: Computation of a tropical rank of a matrix, P3: Computation of optimal strategies in parity games (typically: Mean Payoff Games), P4: Scheduling with and/or precedence constraints, P5: Shortest path problem in hypergraph, P6: Model checking and $\mu$-calculus.
@@ A.2 Choosing a variable for a new substitution after $k$ substitutions in MPR
Because when $\mathbbm{x}^{+}\neq\emptyset$ the cost function can increases we have the following result.
@@
The function $h\mapsto c_{h}\otimes h$ is the lower bound of the cost function $(x,h)\mapsto c^{\intercal}_{\mathbbm{x}}\otimes x\oplus c_{h}\otimes h$ which is attained at $x=\boldsymbol{-\infty}$ iff the following conditions hold:
the cost function cannot increase, ie.:
$\mathbbm{x}^{+}=\emptyset,$ (109a)
and the homogenization variable $h$ can take any arbitrary value, ie.:
$A_{[|1,m|]h}^{+[k]}\leq A_{[|1,m|]h}^{-[k]}.$ (109b)
@@
Case 0: $\max(z)=\mbox{{\scriptsize$\mathbb{O}$}}$ or no maximum. Case $0.1$: $\mathbbm{x}=\emptyset$ and the condition (109b) of Theorem A.2 does not holds. $\mathbbm{x}=\emptyset$ means that $k=n$ (no more variable to be substituted) and $\textsf{cost}^{[n]}(x,h)=c_{h}^{[n]}\otimes h$. The condition (109b) of Theorem A.2 does not holds means $A_{[|1,m|]h}^{+[n]}\nleq A_{[|1,m|]h}^{-[n]}$ then $h=\mbox{{\scriptsize$\mathbb{O}$}}$ is the only possible value for $h$ and by backward substitution we have $z=\mbox{{\scriptsize$\mathbb{O}$}}$, $x=\boldsymbol{-\infty}$.
Case $0.2$: $\mathbbm{x}\neq\emptyset$ and $I_{\leq}(\mathbbm{x})=\emptyset$ and $I_{\geq}(\mathbbm{x})=\emptyset$ and the conditions (109a)-(109b) of Theorem A.2 do not hold. The substitution process stops because no valid inequalities can be generated. No maximum.
Case $1$: switching case, cf. § 2.4. $\mathbbm{x}^{+}=\emptyset$ and $\textsf{cost}^{[k]}(x,h)=c_{h}^{[k]}\otimes h$ and the condition (109b) of Theorem A.2 does not holds and $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}\neq\emptyset$ and $I_{\leq}(\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}})=\emptyset$ and $I_{\geq}(\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}})\neq\emptyset$. The substitution must switch to the following minimizing problem:
@@
The other case is Case $2.2$: $\mathbbm{x}\neq\emptyset$ and $\exists x_{j}\in\mathbbm{x}$: $I_{\leq}(x_{j})\neq\emptyset$. Let us distinguish the following subcases which are based on the form of the cost function:
@@
Case $2.2.2$: otherwise, the conditions (109a)-(109b) of Theorem A.2 do not hold. Based on the hierarchy $\preceq_{\overline{var}}$ (see (105)) between the remaining variables let us define the set of the dominating variables $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}:=\{x_{j}\in\mathbbm{x}:\;I_{\geq}(x_{j})\neq\emptyset\mbox{ and }I_{\leq}(x_{j})\neq\emptyset\}$. And the set of inequalities:
@@
Where: $I_{\leq}(\mathbbm{x}\cap D^{+\mbox{{\scriptsize$\mathbb{O}$}}}):=\{(i,j):\mbox{ $x_{j}\in\mathbbm{x}\cap D^{+\mbox{{\scriptsize$\mathbb{O}$}}}$ and $I_{\leq}(x_{j},i)$ is valid}\}$.
@@ A.3 The new characteristic elements of the MPR after substitution
$\forall j:t^{k\rightarrow k+1}_{j,.}:=\left\{\begin{array}[]{lc}e^{\intercal}_{j}\mbox{ if $j\neq j^{*}$}\\ (\mbox{{\scriptsize$\mathbb{O}$}},v^{\intercal}_{i^{*}j^{*}},r_{i^{*}j^{*}})&\mbox{ if $j=j^{*}$.}\end{array}\right$. (124)
Where $e^{\intercal}_{j}$ denotes the $j$th $n+2$-dimensional row vector of the $(n+2)\times(n+2)$-identity matrix $\boldsymbol{I}_{n+2}$.
@@ A.4 A numerical example
Applying results of § A.3 the new characteristic elements are listed below.
@@
$\mathbbm{x}=\{x_{1}\}\neq\emptyset$. It is easy to see that the MPR problem is bounded. We have $\mathbbm{x}^{+}=\{x_{1}\}$ and $\mathbbm{x}^{\mbox{{\scriptsize$\mathbb{O}$}}}=\emptyset$. The set of dominating variables is $D^{+\mbox{{\scriptsize$\mathbb{O}$}}}:=\{x_{1}\}$. The conditions (109a)-(109b) of Theorem A.2 are not verified. Thus, Case $2.2.2$ applies. We have (see (121)): $\mathbb{I}^{+\mbox{{\scriptsize$\mathbb{O}$}}}(D)=I_{\leq}(\{x_{1}\})=\{(2,1),(3,1),(4,1)\}$.
@@
Applying results of § A.3 the new characteristic elments of the problem are listed below.
@@
Clearly, $A_{[|1,4|]h}^{+[2]}=\boldsymbol{-\infty}\leq A_{[|1,4|]h}^{-[2]}=\boldsymbol{-\infty}$. And the conditions (109a)-(109b) of Theorem A.2 are verified and the cost function $h\mapsto 5\otimes h$ is reached.
The solution of the MPR is obtained by a trivial backward substitution: